The area of a square is the same as its perimeter. The measure of a side is (x – 4)units. What is the only possible value of x?

Answer by CharlesG2(834)

The area of a square is the same as its perimeter. The measure of a side is (x – 4)units. What is the only possible value of x?
area square = side^2 = a^2
perimeter square = 4*side = 4a
a^2 = 4a
a = 4
plug in (x-4) for a
x - 4 = 4
x = 8
(x-4)^2 = 4^2 = 16
4(x-4) = 4x - 16 = 32 - 16 = 16